OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!~~~A car covers 72 kilometers in the first hour of its journey. In the next hour, it covers 90 kilometers. What is the amount of work done by the car? The total mass of the car, including its passengers, is 2.5 × 103 kilograms. 2.8 × 105 joules 4.0 × 105 joules 5.1 × 105 joules 2.8 × 106 joules 2.8 × 107 joules

OpenStudy (loser66):

my calculator is out of battery so that I can't compute anything I give you the step 1/ find the accelerate of the car: the first hour drive 72km means \(\dfrac{72*10^3m}{3600seconds}\) is the velocity v1of the first hour do the same with the second hour to get V2 then V2-V1 = a the total distance it travel is d = 72*10^3 +92*10^3 m apply all to W = F*d = m*a*d where m is mass of the car, a and d are calculated above,

OpenStudy (anonymous):

ummm yeah thats way too confusing @Loser66

OpenStudy (loser66):

How can I make it simpler? It is the only 1 way to do

OpenStudy (anonymous):

I don't know. @Loser66